The diameter of a sphere is measured to be 5.36 in. Find (a) the radius of the sphere in centimeters, (b) the surface area of the sphere in square centimeters, and (c) the volume of the sphere in cubic centimeters.


1
Expert's answer
2021-06-22T09:47:32-0400

(a) Let's first find the radius of the sphere:


"r=\dfrac{D}{2}=\dfrac{5.36\ in\cdot\dfrac{2.54\ cm}{1\ in}}{2}=6.8\ cm."

(b) We can find the surface area of the sphere as follows:


"A=4\pi r^2=4\pi(6.8\ cm)^2=581\ cm^2."

(c) We can find the volume of the sphere as follows:


"V=\dfrac{4}{3}\pi r^3=\dfrac{4}{3}\pi(6.8\ cm)^3=1317\ cm^3."